40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
Message ID: 17042     Entry time: Thu Jul 28 14:34:40 2022     In reply to: 17040     Reply to this: 17044
Author: Yehonathan 
Type: Update 
Category: BHD 
Subject: LO beam power at BHD DCPDs is significantly lower than expected 

{Yuta, Yehonathan}

We went to the BS table to check the POP beam power. We first notice that the POP beam has a nice gaussan profile on the viewing card. We traced it the beam to the viewing port and measured the power there. Before measuring the power we misalign ITMY/ITMX to get rid of interferences. We measure the beam to be 205uW in both cases.

The expected power is

950 mW * 0.9 (IMC transmission?) * 5.637%(PRM) * 97.8%(PR2) * 50%(BS) * 98.6%(ITM) * 50%(BS)  * 2.2%(PR2) = 260uW

which is reasonably close to what we measure which confirms that PR2 transmission is around what we think it is.

This strengthen our suspicion that LO beam gets clipped somewhere.

 

We also improved the clipping on the POP camera by one of the beamsplitters along the beam path and the alignment to the POPDC PD (~100 cts before, ~ 1000 cts after).

 

ELOG V3.1.3-